OpenStudy (anonymous):

How do I get the x and y intercept of 4-2^x+3?

OpenStudy (anonymous):

The x intercept occurs when y = 0. And the y intercept occurs when x = 0. Since your equation is y = 4 - 2^x +3, simply plug in 0 for one of the two variables, and solve for the other.

OpenStudy (anonymous):

How would I go on solving for x? That's what confuses me. I understand how to get the y.

OpenStudy (anonymous):

0=4-2^x 2^x=4 x = 2 (since 2² = 4)

OpenStudy (anonymous):

Is the equation: \[y=4-2^x + 3\] or \[y=4-2^{x + 3}\]

OpenStudy (anonymous):

oops sorry forgot about the 3

OpenStudy (anonymous):

i need to sleep lol ... sorry

OpenStudy (anonymous):

To get a variable out of the exponent one would typically use a log function. Unless you can do it easily by inspection.

OpenStudy (anonymous):

The second one Polpax. Thanks for trying kiwi :)

OpenStudy (anonymous):

Ok, so \[y=0 \rightarrow 2^{x+3} = 4\] Since we know \[4 = 2^2 \rightarrow 2^{x+3} = 2^{2} \rightarrow x+3 = 2 \rightarrow x = ?\]

OpenStudy (anonymous):

Thanks Polpak :)